A Country’s Government has approved a relocation plan for a section of its population to inhabit a 3-island portion. Design a hypothetical Telecommunication network(Fixed and mobility consideration) for both internal and external communication.

Consider the following

Islands location is 5000km from the main land 

Islands are equidistant from one another

Accessibility by water transport

Government very resourceful – enough money

A lot of economic activity on the Islands (billions of dollars annually)

1.5 million population with 80% adults

120-kilometre bridge links the three islands


1
Expert's answer
2021-03-12T05:42:33-0500

Given,

Distance between Iceland location and main Iceland = 5000km

Total population = 1.5 million

Percentage of adults = 80%

bridge link = 120km

"d=\sqrt{2Rh}"

"=\sqrt{2\times 120\times 5000}km"

"=1095.44km"